# Searching for Exotrojans in Pulsar Binary Systems

### v1

@@ 3.1 Conjunction Inconsistency
The reflex motion of the pulsar due to the orbit of its companions is fully consistent with the replacement of the dominant companion and exotrojan with a single effective object (A. Leleu et al. 2017; see Figure 1). If the time of the dominant companion’s furthest (closest) approach to Earth—known as superior (inferior) conjunction—can be measured directly through optical band data, then this time, $T_{\rm o,s}$ ($T_{\rm o,i}$), can be compared to the conjunction time as predicted by the pulsar’s reflex motion, $T_{\rm r,s}$ ($T_{\rm r,i}$). Given how C. Lange et al. (2001) define $T_{\rm asc}$, radio pulsar timing predicts superior and inferior conjunctions to occur at $T_{\rm r,s}=T_{\rm asc}-P_{\rm b}/4$ and $T_{\rm r,i}=T_{\rm asc}+P_{\rm b}/4$, respectively, where we ignore $\mathcal{O}(e_{\rm o})$ effects. $P_{\rm b}$ is the binary orbital period and $T_{\rm asc}$ is the time of the ascending node as defined in C. Lange et al. (2001). E. B. Ford & B. S. Gaudi (2006) show that any discrepancy in $T_{\rm o,s}$ and $T_{\rm r,s}$ (or $T_{\rm o,i}$ and $T_{\rm r,i}$) could indicate the presence of an exotrojan with the time difference directly mapped from the exotrojan mass:
$\Delta t\approx\pm 35.6\left(\frac{P}{3\>\text{hours}}\right)\left(\frac{0.04M_{\odot}}{m_{\rm T}+m_{\rm c}}\right)\left(\frac{\Delta m_{\rm T}}{1\>M_{\rm J}}\right)\>\text{seconds},$ (2)
@@
where $\Delta t=T_{\rm o,s}-T_{\rm r,s}$ and $\Delta m_{\rm T}$ is the mass difference between the $L_{4}$ and $L_{5}$ Lagrange points. Here, we ignore the $\mathcal{O}(e_{\rm o})$ term. Therefore, Trojans present in both points will partially cancel their measurable effects. We use this method to rule out a high-mass Trojan in the PSR J1641+8049 system (see Section 4.1) by incorporating its superior conjunction measurements (Section 2.1). We note that this method is inclination independent due to the orbital period and the time delay being inclination independent.
Figure 1: A schematic of the positions of the Trojan, $m_{\rm T}$, the dominant companion, $m_{\rm c}$, and the effective combined object, $m_{\rm f}$. $\Delta\phi$ is the relative angle between the orbital longitudes of the combined companion and the actual dominant companion. The Trojan and dominant companion are shown orbiting counterclockwise here, meaning the Trojan is in the $L_{4}$ Lagrange point of the pulsar-dominant companion system. At the time of this figure, the companion is at its closest position to Earth, known as inferior conjunction. The line of sight to the Earth is shown in red. See the text for more details. Inspired by Figure 1 of E. B. Ford & B. S. Gaudi (2006).
3.1.1 Determination of the Times of Optical Superior Conjunction
@@
The PSR J1641+8049 optical light curves show a clearly periodic structure with a period equal to the binary period (e.g., Figure 4 of A. Y. Kirichenko et al. 2024), explained by the neutron star heating its tidally locked companion. Therefore, assuming the symmetric heating models of S. Zharikov et al. (2013, 2019) and neglecting the heating of the Trojan, the peak light corresponds to the time(s) of superior conjunction of the dominant companion.
@@
To derive the times of the superior conjunction and their associated uncertainties, we adopted the HiPERCAM dataset due to its superior time resolution. We show the optical light curve in Figure 3. More details can be found in D. Mata Sánchez et al. (2023); A. Y. Kirichenko et al. (2024). We adopt the emission model of S. Zharikov et al. (2013, 2019) and start with the best-fit parameters found in A. Y. Kirichenko et al. (2024). These parameters are the distance from the system to the Earth, the reddening E(BV)E(B-V), the pulsar mass, the companion-to-pulsar mass ratio, the orbital inclination, the Roche lobe filling factor, the effective radiation factor, the companion dark-side temperature, and the time of superior conjunction. For more details on these parameters and the model, see S. Zharikov et al. (2013, 2019); A. Y. Kirichenko et al. (2024) and references therein. We define $\mathbf{p}$ as the vector of all these parameters, excluding the time of superior conjunction.
@@
3.2 Librations
A. Leleu et al. (2015) adapted the work of P. Robutel et al. (2011) to determine how a host star’s position (or velocity if spectral data is available) relative to the Solar System barycenter (SSB) will change when a companion and Trojan are present. A pulsar moving relative to the SSB will cause changing light travel time delays, known as Rømer delays, in the pulse TOAs. R. J. Jennings et al. (2020b), for example, showed these delays to be
@@ 3.2 Librations
when ignoring terms that are degenerate with the pulsar’s pulse period and pulse period time derivative. $\boldsymbol{r}_{\rm p}$ is the pulsar’s position with respect to the system center of mass, $\boldsymbol{\hat{n}}$ is the unit vector pointing from the pulsar system barycenter to the SSB, and $c$ is the speed of light in a vacuum. Using Equation 4 with the results of A. Leleu et al. (2015), we find (Appendix A)
$+\frac{2}{3}\frac{\nu}{n_{\rm b}}\sin{(\nu\tilde{t}+\varphi)}\sin{(n_{\rm b}\tilde{t}+\theta)}\Big]$ (5)
in the small-libration limit, where $\nu$ is the libration (angular) frequency, $\varphi$ is the arbitrary phase of the Trojan at $t=T_{\rm asc}$, $\theta$ is a phase that depends on $\delta=m_{\rm c}/(m_{\rm c}+m_{\rm T})$ (see Equation 9), $t$ is time, $\tilde{t}\equiv t-T_{\rm{asc}}$, $n_{\rm{b}}$ is the orbital mean motion (angular) frequency, and $\mathcal{T}$ is the libration strength (which has units of time). Equation A19 is an alternative and equivalent representation of Equation 3.2 where we cast the signal in terms of its fundamental frequencies. We choose to make $\tilde{t}$ relative to $T_{\rm{asc}}$ because a Fisher information analysis [^3] We point to D. Wittman (2025) for an excellent overview of the Fisher information matrix for beginners. reveals that the uncertainties in $\theta$ and particularly $\varphi$ are minimized for a central reference epoch with respect to the TOAs. Furthermore, picking $T_{\rm{asc}}$ rather than the TOA average constrains $\theta$ to directly reveal $\delta$ (see Appendix A). Although $\nu$ is an angular frequency, we report physically determined values in ordinary frequency units. The low-libration limit assumes the relative longitudes of the companion and Trojan, $\zeta$, can be modeled as
$\zeta=z\cos{(\nu t+\varphi)}+\frac{\pi}{3},$ (6)
where $z$ is the libration amplitude. See Figure 2 and Appendix A for the derivation and more details.
[…]
$\mathcal{T}$ $=\alpha z\delta(1-\delta),$ (7)
$K$ $=\alpha\sqrt{1-\delta(1-\delta)},$ (8)
$\alpha=\mu a\sin{(I)}/c$, $\mu=(m_{\rm T}+m_{\rm c})/(m_{\rm p}+m_{\rm c}+m_{\rm T})$, and $a$ is the mass-averaged semi-major axis of the companion and Trojan in the pulsar-centric reference frame. Additionally, A. Leleu et al. (2015) showed that in the low $\mu$ limit.
$\nu=n_{\rm b}\sqrt{\frac{27\mu}{4}}.$ (10)
@@
where $\kappa=(m_{\rm p}-m_{\rm c})/(m_{\rm p}+m_{\rm c})$ (N. J. Cornish, 2001; B. Érdi et al., 2007). Equation 11 is more appropriate than Equation 10 when $m_{\rm T}$ is negligible but $m_{\rm c}$ is not. There is another libration frequency as well (B. Érdi et al., 2007), but M. Janson (2013) showed that its corresponding libration amplitude is much smaller than $z$, and therefore can be safely ignored when $\mathcal{T}$ is already near the threshold of detectability, as is the case for the pulsar systems in our study (Section 4.2).
We solve Equations 7, 8, and 9 in reverse order analytically:
@@
Upon assuming a pulsar mass, the companion and Trojan masses are
@@
The orbital inclination can be determined too, using the Newtonian binary mass function as formulated in P. C. Freire & N. Wex (2024),
@@
where $T_{\odot}=4.925490947641...\,\rm{\mu s}$ is the nominal solar mass parameter. [^4] This is a defined quantity. See Section 3.4.1 of P. C. Freire & N. Wex (2024) for a brief discussion on $T_{\odot}$.
Figure 2: The relevant angles in the triangular three-body system when librations are present. The Earth lies in the xzxz-plane, where $\hat{z}=\hat{x}\times\hat{y}$, with the line of sight making an angle $I$ with the $z$-axis. $m_{\rm f}$ is the fictitious companion mass given by Equation A14. The Trojan and companion are shown orbiting counterclockwise here, meaning the Trojan is in the $L_{4}$ Lagrange point. If $m_{\rm T}>m_{\rm c}$, then the companion and Trojan should be interchanged by our convention. $\lambda$ is the angle that the Trojan and dominant companion librate around, which can differ slightly from $\phi$, the angle given by the location of the combined fictitious companion. See the text for more details. Inspired by Figure 1 of A. Leleu et al. (2015).
3.2.1 Determination of Libration Parameters
@@
We adapt the search methodology of L. Dey et al. (2025), originally developed for freely floating interstellar objects (FFOs), to our Trojan search. In order to seek out the libration delays of Equation 3.2, we model not only $\Delta\tau$, but also white noise, red noise, and the ordinary pulsar timing model, which includes parameters such as pulsar spin period, spin period derivatives, sky position, and binary parameters (see R. T. Edwards et al. 2006; J. Luo et al. 2019, 2021; A. Susobhanan et al. 2024 for more-detailed discussions on the ordinary pulsar timing model). We model the noise and timing model in the exact same way as L. Dey et al. (2025). Noise and timing modeling are handled in the ENTERPRISE Python library (J. A. Ellis et al., 2019; A. D. Johnson et al., 2023). We also use the PTMCMCSampler Python library (J. Ellis & R. van Haasteren, 2017; A. D. Johnson et al., 2023), which is an implementation of the parallel tempering Markov-Chain Monte Carlo (PTMCMC) method (see, for example, D. J. Earl & M. W. Deem 2005; W. Vousden et al. 2016 for discussions on PTMCMC). Unlike L. Dey et al. (2025), we used sixteen temperature chains rather than just one because the posterior space in our case can be highly multimodal (see Section 4.3).
@@
We do not have to worry about degeneracies between the noise modeling and the libration frequency. The red noise is modeled as frequencies in integer multiples of $1/T_{\rm span}$, where $T_{\rm span}$ is the time span of TOAs for a given pulsar, up to $30/T_{\rm span}$. The highest such frequency for the eight NG15yr pulsars of interest is $30/(3.4\,{\rm yr})\sim 9\,\rm{yr}^{-1}$. The libration frequencies are of order $1\,\rm{day}^{-1}\sim 400\,\rm{yr}^{-1}$ (see Figures 4, 5, and 6). Therefore, the red noise frequency space and the libration frequency space are disjoint.
3.2.2 Libration Priors

### v2

@@ II.1 Conjunction Inconsistency
Let $T_{\rm o,s}$ and $T_{\rm o,i}$ be the times of superior and inferior conjunction, respectively, as measured through optical band data, and let $T_{\rm r,s}$ and $T_{\rm r,i}$ be the same as measured by the TOA-determined pulsar reflex motion. Given how Lange et al. (2001) define $T_{\rm asc}$, radio pulsar timing predicts superior and inferior conjunctions to occur at $T_{\rm r,s}=T_{\rm asc}-P_{\rm b}/4$ and $T_{\rm r,i}=T_{\rm asc}+P_{\rm b}/4$, respectively, where we ignore orbital eccentricity ($e_{\rm o}$) effects. $P_{\rm b}$ is the binary orbital period and $T_{\rm asc}$ is the time of the ascending node as defined in Lange et al. (2001). We ignore $\mathcal{O}(e_{\rm o})$ effects throughout this study, because the highest orbital eccentricity of the pulsars studied is of the order $10^{-4}$. Ford and Gaudi (2006) show that any discrepancy in $T_{\rm o,s}$ and $T_{\rm r,s}$ (or $T_{\rm o,i}$ and $T_{\rm r,i}$) could indicate the presence of an exotrojan with the time difference directly mapped from the exotrojan mass:
$\Delta t\approx\pm 35.6\left(\frac{P_{\rm b}}{3\,\text{h}}\right)\left(\frac{0.04\,{\rm M}_{\odot}}{m_{\rm T}+m_{\rm c}}\right)\left(\frac{\Delta m_{\rm T}}{1\,\rm{M}_{\rm J}}\right)\,\text{s},$ (2)
@@
where $\Delta t=T_{\rm o,s}-T_{\rm r,s}$ and $\Delta m_{\rm T}$ is the mass difference between the $L_{4}$ and $L_{5}$ Lagrange points. Here, again, we ignore the $\mathcal{O}(e_{\rm o})$ term. Therefore, Trojans present in both points will partially cancel their measurable effects. We use this method to rule out a high-mass Trojan in the PSR J1641+8049 system (see Section IV.1) by incorporating its superior conjunction measurements (Section III.1). We note that this method is inclination independent due to the orbital period and the time delay both being inclination independent.
II.1.1 Determination of the Times of Optical Superior Conjunction
@@
Some binary pulsar optical light curves show a clearly periodic structure with a period equal to the binary period (e.g., Figure 4 of Kirichenko et al. 2024), explained by the neutron star heating its tidally locked companion. Therefore, assuming the symmetric heating models of Zharikov et al. (2013, 2019) and neglecting the heating of the Trojan, the peak flux corresponds to the time(s) of superior conjunction of the secondary.
@@
To derive the times of the superior conjunction and their associated uncertainties, we analyzed light curve data of the PSR J1641+8049 system. We show the optical light curve in Figure 1. More details can be found in Section III.1 and Mata Sánchez et al. (2023); Kirichenko et al. (2024). We adopt the emission model of Zharikov et al. (2013, 2019) and start with the best-fit parameters found in Kirichenko et al. (2024). The emission model assumes symmetric heating with hottest side facing the pulsar. The heating model parameters are the distance from the system to the Earth, the reddening E(BV)E(B-V), the pulsar mass, the companion-to-pulsar mass ratio, the orbital inclination, the Roche lobe filling factor, the effective radiation factor, the companion dark-side temperature, and the time of superior conjunction. For more details on these parameters and the model, see Zharikov et al. (2013, 2019); Kirichenko et al. (2024) and references therein. We define $\mathbf{p}$ as the vector of all these parameters, excluding the time of superior conjunction.
Figure 1: High PERformance CAMera light curve data from the 10.4m Gran Telescopio Canarias of the PSR J1641+8049 system for each Sloan band. The data was first presented in Mata Sánchez et al. (2023). We show the best-fit model for the $r^{\prime}$ band as an example, with the remaining model curves shown in Kirichenko et al. (2024). Every band except $u^{\prime}$ was used in calculating $T_{\rm o,s}$. The vertical dashed line marks the TSC as predicted by radio timing. The uncertainty on the TOA-predicted TSC is far smaller than the $\hat{x}-$axis scale of the figure. The vertical cyan band encloses the 68% confidence interval of the TSC as calculated from the optical light curve.
@@
II.2 Librations
Leleu et al. (2015) adapted the work of Robutel et al. (2011) to determine how a host star’s position (or velocity if spectral data is available) relative to the Solar System barycenter (SSB) will change when a companion and Trojan are present. A pulsar moving relative to the SSB will cause changing light travel time delays, known as Rømer delays, in the pulse TOAs. Jennings et al. (2020a), for example, showed these delays to be
@@ II.2 Librations
when ignoring terms that are degenerate with the pulsar’s pulse period and pulse period time derivative. $\boldsymbol{r}_{\rm p}$ is the pulsar’s position with respect to the system center of mass, $\boldsymbol{\hat{n}}$ is the unit vector pointing from the pulsar system barycenter to the SSB, and $c$ is the speed of light in a vacuum. Therefore, $\boldsymbol{\hat{n}}$ is antiparallel to the LOS.
We will now begin the derivation of $\Delta\tau$ in the case of a librating Trojan in the orbit around a pulsar, but reserve most of the details to Appendix A. Let $\lambda_{\rm T}$ and $\lambda_{\rm c}$ be the orbital longitudes of the Trojan and secondary, respectively. Let $\lambda$ be the mass-averaged longitude of the co-orbitals which follows the mean motion and is unaffected by librations. We call the difference in longitude between the co-orbitals the libration angle, $\zeta$. The TOAs will indicate the presence of an orbiting combined effective companion, which has longitude $\phi$. $\lambda$ and $\phi$ are nearly but not exactly identical, with
$|\phi-\lambda|\lesssim 0.02$ (5)
shown in Appendix A. Both $\lambda$ and $\phi$ evolve over time as the co-orbitals orbit the pulsar according to
$\lambda=n_{\rm b}\tilde{t}+\lambda_{0},$ (6)
$\phi=n_{\rm b}\tilde{t}+\phi_{0},$ (7)
where $n_{\rm b}=2\pi/P_{\rm b}$ is the orbital mean motion, $\tilde{t}$ is time with respect to the time of the ascending node, $T_{\rm asc}$, and $\lambda_{0}$ and $\phi_{0}$ are to-be-determined constants. We show in Appendix A that the mass of the fictitious effective companion
$m_{f}=(m_{\rm c}+m_{\rm T})\sqrt{1-\delta(1-\delta)}+\mathcal{O}\left(\frac{m_{\rm T}}{m_{\rm p}},\,\frac{m_{\rm c}}{m_{\rm p}}\right).$ (8)
Defining
$\delta=m_{\rm c}/(m_{\rm c}+m_{\rm T}),$ (9)
the mass of the secondary relative to both co-orbitals, the angles of the Trojan and secondary longitudes with respect to $\lambda$ are $\delta\zeta$ and $(1-\delta)\zeta$, respectively. Robutel et al. (2011) showed that
$\lambda_{\rm T}$ $=\lambda+\delta\zeta,$ (10)
$\lambda_{\rm c}$ $=\lambda-(1-\delta)\zeta.$ (11)
Let the plane of the orbit define the $\hat{x}\hat{y}-$plane. Without a loss of generality, we place the Earth in the $\hat{x}\hat{z}-$plane and define the $\hat{x}-$axis to be the projection of the LOS onto the orbital plane. Therefore, the orbital inclination with respect to the plane of the sky, $I$, is the angle between the LOS and the $\hat{z}-$axis. We define the $\hat{y}-$axis such that the $T_{\rm asc}$ is the time at which the fictitious companion crosses the negative $\hat{y}-$axis. See Figure 2 for a helpful illustration.
Figure 2: The relevant angles in the triangular three-body system. $m_{\rm f}$ indicates the fictitious companion, with mass given by Equation 8. The Trojan and companion are shown orbiting counterclockwise here, meaning the Trojan is in the $L_{4}$ Lagrange point. See the text for more details. Inspired by Figure 1 of Leleu et al. (2015).
When librations are absent, $\zeta=\pi/3$ for all time. However, in the presence of small librations, $\zeta$ will oscillate around $\pi/3$ with (angular) frequency $\nu$ and amplitude $z$ (e.g., Leleu et al., 2015); i.e.
$\zeta=z\cos{(\nu\tilde{t}+\varphi)}+\frac{\pi}{3},$ (12)
where $\varphi$ is the vibrational phase at $t=T_{\rm asc}$. $z$ can vary across otherwise identical systems and is a function of the Trojan’s detailed formation history. We note that $z$ can be lowered through any sort of dissipative process, such as gas drag (Lyra et al., 2009). In contrast, $\nu$ depends on the masses of the pulsar and co-orbitals. Namely, Leleu et al. (2015) showed that in the low $\mu$ limit,
$\nu=n_{\rm b}\sqrt{\frac{27\mu}{4}},$ (13)
where
$\mu=\frac{m_{\rm T}+m_{\rm c}}{m_{\rm p}+m_{\rm c}+m_{\rm T}}.$ (14)
@@
where $\kappa=(m_{\rm p}-m_{\rm c})/(m_{\rm p}+m_{\rm c})$ (Cornish, 2001; Érdi et al., 2007). Equation 15 is more appropriate than Equation 13 when $m_{\rm T}$ is negligible but $m_{\rm c}$ is not. There is another frequency associated with librations as well (Érdi et al., 2007), which would add in sum to Equation 12, but Janson (2013) showed that its corresponding libration amplitude is much smaller than $z$, and therefore can be safely ignored when these systems are on the threshold of detectability.
If $z=0$, then the TOAs are completely degenerate from TOAs from a two-body system. A non-zero $z$ breaks this degeneracy. The TOA signal, through Equation 4, will have a term degenerate with a two-body system plus a term due to librations. This “libration term” will depend on the mass-averaged semi-major axis of the co-orbitals in the pulsar-centric frame ($a$), $z$, $\nu$, $n_{\rm b}$, $I$, and $\mu$. After removing the degenerate two-body term, we derive in Appendix A
$\Delta\tau=$ $\frac{\mu a\sin{I}}{c}z\delta(1-\delta)\Big[\cos{(\nu\tilde{t}+\varphi)}\cos{(n_{\rm b}\tilde{t}+\theta)}$
$+\frac{2}{3}\frac{\nu}{n_{\rm b}}\sin{(\nu\tilde{t}+\varphi)}\sin{(n_{\rm b}\tilde{t}+\theta)}\Big],$ (16)
where
In searching for a Trojan in a particular system, see Section II.2.1, we wish to remove covariances and degeneracies in the prefactor of Equation II.2, so we define what we call the libration strength
$\mathcal{T}=\alpha z\delta(1-\delta),$ (18)
where
$\alpha=\frac{\mu a}{c}\sin{(I)}.$ (19)
Therefore, we have
$+\frac{2}{3}\frac{\nu}{n_{\rm b}}\sin{(\nu\tilde{t}+\varphi)}\sin{(n_{\rm b}\tilde{t}+\theta)}\Big].$ (20)
We plot this function for an example set of parameters in Figure 3. The timing model, Equation 4, would cause TOA residuals, which are the measured TOAs minus the predicted TOAs, if a Trojan in the system was not accounted for.
[…]
While pulsar timing cannot reveal the semi-major axis of the pulsar’s orbit around the system center of mass ($a_{\rm p}$) directly, it can measure its projection onto the plane of the sky
$K\equiv\frac{a_{\rm p}}{c}\sin{I}.$ (21)
This is useful quantity because it can be related to the masses of the co-orbitals through
$K=\alpha\sqrt{1-\delta(1-\delta)}$ (22)
by setting $K$ to be the amplitude of the two-body degenerate term of the three-body system (see Appendix A). The value of $K$ is known quantity for any timed binary pulsar a priori to any Trojan search.
[…]
@@
Therefore, Equation 23 shows that $\theta$ gives a direct measure of the relative Trojan mass, $\delta$. Upon assuming a pulsar mass, the companion and Trojan masses are
@@
The orbital inclination can be determined too, using the Newtonian binary mass function
@@
where $T_{\odot}=4.925490947641...\,\rm{\mu s}$ is the nominal solar mass parameter. [^2] This is a defined quantity. See Section 3.4.1 of Freire and Wex (2024) for a brief discussion on $T_{\odot}$.
II.2.1 Determination of Libration Parameters
@@
We adapt the search methodology of Dey et al. (2026), originally developed for freely floating interstellar objects (FFOs), to our Trojan search. In order to seek out the libration delays of Equation II.2, we model not only $\Delta\tau$, but also white noise, red noise, and the ordinary pulsar timing model, which includes parameters such as pulsar spin period, spin period derivatives, sky position, and binary parameters (see Edwards et al. 2006; Luo et al. 2019, 2021; Susobhanan et al. 2024 for more-detailed discussions on the ordinary pulsar timing model). We model the noise and timing model in the exact same way as Dey et al. (2026). Noise and timing modeling are handled in the ENTERPRISE Python library (Ellis et al., 2019; Johnson et al., 2023). We also use the PTMCMCSampler Python library (Ellis and van Haasteren, 2017; Johnson et al., 2023), which is an implementation of the parallel tempering Markov-Chain Monte Carlo (PTMCMC) method (see, for example, Earl and Deem 2005; Vousden et al. 2016 for discussions on PTMCMC). Unlike Dey et al. (2026), we used sixteen temperature chains rather than just one because the posterior space in our case can be highly multimodal (see Section IV.3).
@@
We do not have to worry about degeneracies between the noise modeling and the libration frequency. The red noise is modeled as frequencies in integer multiples of $1/T_{\rm span}$, where $T_{\rm span}$ is the time span of TOAs for a given pulsar, up to $30/T_{\rm span}$. The highest such frequency for the eight NG15yr pulsars of interest is $30/(3.4\,{\rm yr})\sim 9\,\rm{yr}^{-1}$. The libration frequencies are of order $1\,\rm{day}^{-1}\sim 400\,\rm{yr}^{-1}$ (see Figures 5, 6, and 7). Therefore, the red noise frequency space and the libration frequency space are disjoint.
II.2.2 Libration Priors